Kirkland is a small country described by the following, Y =8000;G=2500;T =2000;
C = 1000 + .66(Y − T )
I = 1200 − 100r
Y =C+I+G
(a) For Kirkland, compute private saving, public saving and national
saving?
(b) Find the equilibrium interest rate?
1
(c) Suppose G is reduced by 500, compute private saving, public sav- ing and national saving?
(d) Find the new equilibrium interest rate?
1
Expert's answer
2020-05-08T14:43:32-0400

(a) For Kirkland, compute private saving, public saving and national

saving?

"\text{Public saving = T - G}"

"\text{Public saving} = 2000-2500=-500"


"\text{Private saving = Y -T -C}"

"\text{Private saving}=8,000 - 2,000 - (1000 + 0.66 * (8,000-2000)"

"\text{Private saving}=1040"


"\text{National savings = Private savings + Public savings}"

"\text{National savings}=1040-500=540"


(b) Find the equilibrium interest rate?

The equilibrium position stipulates that:

"\text{ Equilibrium, Y = C+ I + G}"

"8,000 = (1000 + 0.66 * (Y-T)) + 2,500 + 1,200 - 100r"

"8,000 = (1000 + 0.66 * (8000-2000) + 2,500 + 1,200 - 100r"

"100r=8660"


"r=\dfrac{8660}{100}=8.66"


(c) Suppose G is reduced by 500, compute private saving, public saving and national saving?

By reducing G by 500 the new G will be 200.

The public investment will be:

"\text{Public saving = T - G}=2000-2000=0"

The private investment will be:

"\text{Private saving}=8,000 - 2,000 - (1000 + 0.66 * (8,000-2000)"


The national savings will therefore, be:

"\text{National savings}=1040-0=1040"


(d) Find the new equilibrium interest rate?

"8,000 = (1000 + 0.66 * (Y-T)) + 2,000 + 1,200 - 100r"

"8,000 = (1000 + 0.66 * (8000-2000) + 2,000 + 1,200 - 100r"

"100r=160"


"r=\dfrac{160}{100}=1.6"